OpenStudy (anonymous):

I need help with integrals. I don't understand the basic rules. Therefore, I cant evaluate. Help would be much appreciated.

OpenStudy (anonymous):

What do you not understand??

OpenStudy (anonymous):

I was absent when my teacher introduced integrals. I don't know how to evaluate.

OpenStudy (anonymous):

For example: \[\int\limits_{1/2}^{3} (2-\frac{ 1 }{ x }) dx\]

OpenStudy (anonymous):

the integral is the anti derivative..so if you have y=f(t) the integral of f'(t)=f(t)..

OpenStudy (anonymous):

for that problem you can integrate is a difference..so this case integral of 2dx - integral 1/xdx

OpenStudy (anonymous):

I thought I was supposed to find the antiderivative of inside the parentheses and evaluate that or something..sigh..idk.

OpenStudy (anonymous):

you need to look up the rules and see how they are defined using the textbook..if you understand derivatives you shouldnt have a problem with these introductory problems

OpenStudy (anonymous):

I have my textbook. I don't understand. That is why I am asking for help. Yet you're telling me to consult my textbook. Gee, thanks.

OpenStudy (anonymous):

did u learn F(b) - F(a) yet? if u did, just do antiderivative of the stuff inside the parenthesis then do F(b) - F(a) for example, \[\int\limits_{a}^{b} f(x)dx\] say F(x) is antiderivative of f(x) then it'd be F(b) - F(a)

OpenStudy (anonymous):

Yes, I did learn F(b)- F(a).

OpenStudy (anonymous):

then just find antiderivative of inside and do F(b) - F(a)

OpenStudy (anonymous):

is the antiderivative 2x-lnx ?

OpenStudy (anonymous):

sorry i haven't gotten to ln's yet. but i can tell u that the 2x part is correct.

OpenStudy (anonymous):

Um, ok. So if you took the antiderivative of 2- (1/x) what would you get?

OpenStudy (anonymous):

Remember that \(\frac{d}{dx} [2x+c] = 2\) and that \(\frac{d}{dx}\ln(x) = 1/x\)

OpenStudy (anonymous):

Right, so I got 2x-ln(x)

OpenStudy (anonymous):

So \(\frac{d}{dx} 2x - \ln (x) + C = 2-1/x\)

OpenStudy (anonymous):

Yes, exactly

OpenStudy (anonymous):

Ok, but now what do I do?

OpenStudy (anonymous):

\[ b = 2 \\ a = 1/2 \\ F(x) = 2x - \ln(x) +C \\ \implies \\ F(b) - F(a) = [2(2) - \ln(2) +C] - [2(1/2) - \ln(1/2) +C] \]

OpenStudy (anonymous):

b=3

OpenStudy (anonymous):

So, [2(3)-ln(3)+C] - [2(1/2)-ln(1/2)+C] ?

OpenStudy (anonymous):

technically, you don't really need to worry about the C's because they'll cancel out.

OpenStudy (anonymous):

and yes. if all your other steps are right.

OpenStudy (anonymous):

Can you help me simplify? I think im doing it wrong

OpenStudy (anonymous):

what did u get?

OpenStudy (anonymous):

umm..[6-ln(3)] - [1-ln(1/2)]

OpenStudy (anonymous):

yeah that's what i got too.

OpenStudy (anonymous):

Is there a way to simplify further?

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=integrate+%282-1%2Fx%29+from+1%2F2+to+3 the log should be a ln

OpenStudy (anonymous):

Ohh, ok, thanks.